OpenStudy (anonymous):

How to prove (1-secx)/(1-cosx)=-secx

OpenStudy (anonymous):

?

OpenStudy (mathstudent55):

Keep in mind: \( \sec \theta = \dfrac{1}{\cos \theta} \)

OpenStudy (anonymous):

I know that, but How would I do the left side.

OpenStudy (mathstudent55):

@Mertsj (1-secx)\( \color{red}{/} \)(1-cosx)=-secx

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

-secx becomes -1/cosx

OpenStudy (mathstudent55):

\( \dfrac{1 - \sec x}{1 - \cos x} = - \sec x\) \( \dfrac{1 - \dfrac{1}{\cos x}}{1 - \cos x} = - \sec x\) \( \dfrac{ \dfrac{\cos x}{\cos x} - \dfrac{1}{\cos x}}{1 - \cos x} = - \sec x\) \( \dfrac{ \dfrac{\cos x - 1}{\cos x} } {1 - \cos x} = - \sec x\) \( \dfrac{ \dfrac{-(1 - \cos x)}{\cos x} } {1 - \cos x} = - \sec x\)

OpenStudy (mertsj):

Sorry. I didn't see that division sign.

OpenStudy (mathstudent55):

\(\dfrac{ {-(1 - \cos x)} } {(1 - \cos x) \cos x} = - \sec x \) \(\dfrac{ -1 } {\cos x} = - \sec x \)

OpenStudy (anonymous):

In step 2, how did you multiply cosx/cosx?

OpenStudy (anonymous):

Wait, never mind

OpenStudy (anonymous):

so its false?

OpenStudy (mathstudent55):

Since 1/cos x = sec x, it is true.